Find g(a+1) when g(x)= 1/4x - 3.
Answer by nycsharkman(136)

Find g(a+1) when g(x)= 1/4x - 3.
Do you see x in the function g(x)?
Replace every x with (a + 1) and simplify.
g(a + 1) = 1/4(a + 1) - 3
g(a + 1) = 1/4a + 4 - 3
g(a + 1) = 1/4a + 1
Did you follow? I just found the value of g(a + 1) and it is the fraction
1/4a + 1